### v2

# A Novel Mathematical Framework for Objective Characterization of Ideas through Vector Embeddings in LLM

Abstract
The demand for innovation in product design necessitates a prolific ideation phase. Conversational AI (CAI) systems that use Large Language Models (LLMs) such as GPT (Generative Pre-trained Transformer) have been shown to be fruitful in augmenting human creativity, providing numerous novel and diverse ideas. Despite the success in ideation quantity, the qualitative assessment of these ideas remains challenging and traditionally reliant on expert human evaluation. This method suffers from limitations such as human judgment errors, bias, and oversight. Addressing this gap, our study introduces a comprehensive mathematical framework for automated analysis to objectively evaluate the plethora of ideas generated by CAI systems and/or humans. This framework is particularly advantageous for novice designers who lack experience in selecting promising ideas. By converting the ideas into higher dimensional vectors and quantitatively measuring the diversity between them using tools such as UMAP, DBSCAN and PCA, the proposed method provides a reliable and objective way of selecting the most promising ideas, thereby enhancing the efficiency of the ideation phase.
keywords:
Ideation; Evaluation; Selection; UMAP; PCA; DBSCAN; LLM; Mathematical Framework; Embedding; Conversational AI
[…]
2_bibliography.bib B. Sankar]sankarb@iisc.ac.in
Introduction
The advent of Conversational AI (CAI) technologies, particularly large language models (LLMs) like GPT, Llama, Gemini, etc., has revolutionized the field of natural language processing (yuhan-etal-2023-unleashing). These tools offer unprecedented capabilities in generating coherent and contextually relevant text, facilitating new forms of co-creation and collaboration in both individual and group settings (meyer2023a). In recent times, since its major outbreak, these models have found their way into several applications, such as Content Generation, Customer Service, Education, Healthcare, Research and Development, Entertainment, Translation and Localization, Personal Assistants, etc. (Desmond_Ashktorab_Pan_Dugan_Johnson_2024). Among these applications, the ability to generate new human-like content holds significant value in various domains. One such field that stands to benefit greatly is product design (Gonzalez_Moran_Houde_He_Ross_Muller_Kunde_Weisz). Particularly during the conceptual design phase, designers are tasked with generating ideas—a process that involves the creation of new content (Liu_Iter_Xu_Wang_Xu_Zhu_2023). Recently, the application of Conversational Artificial Intelligence (CAI) models in this context has garnered increasing attention (Shaer_Cooper_Kun_Mokryn). The authors of this paper have also previously published their research on utilizing CAI as a tool for idea generation in product design.
Conversational AI models have significant implications for the creative content generation process, encompassing both the divergence stage of idea generation and the convergence stage of evaluation and selection of ideasshaer2024aiaugmented. During the divergence stage, LLMs can generate numerous ideas rapidly, providing a rich pool of options for further refinement and selection. This capability is particularly advantageous in brainstorming sessions, where the goal is to produce a wide array of ideas without immediate judgment (Shaer_Cooper_Kun_Mokryn). The increasing availability of LLMs, such as GPT, has enabled widespread adoption in various domains, including creative writing, design, and problem-solving (Shaer_Cooper_Mokryn_Kun_Shoshan_2024). These models can generate text that is often indistinguishable from human-written content, making them valuable tools for augmenting human creativity shaer2024aiaugmented during the ideation process. However, the integration of LLMs into the creative process also necessitates robust evaluation mechanisms to ensure the quality and relevance of the generated ideas.
[…]
A crucial aspect of using LLMs in idea generation is distinguishing between AI-generated and human-generated content. Participants in various studies have emphasized the importance of clearly identifying AI contributions, often using visual cues like icons and outlines gonzalez-a. This distinction not only aids in assigning credit but also helps maintain accountability and ethical considerations in collaborative settings. For instance, capturing the prompt that resulted in the AI output and acknowledging the user who wrote the prompt can provide transparency and accountability in the ideation process.
The ability to generate a diverse range of ideas is vital for effective ideation. Participants have recognized the need for generating unexpected or provocative ideas to stimulate creative thinking gonzalez-a. Novelty and diversity are essential dimensions of idea quality, as they contribute to the originality and uniqueness of the generated ideas (Fiorineschi_Rotini_2023). While traditional tools may lack controls for low-level generative parameters, modern LLMs can be prompted to evaluate their own ideas for relevance, thereby filtering out less pertinent suggestions and enhancing the overall quality of the ideation process. For example, adjusting the temperature parameter in LLMs can control the randomness and creativity of the generated ideas. Higher temperatures result in more diverse and creative outputs, while lower temperatures produce more focused and predictable ideas.
Group brainstorming sessions often face barriers such as peer judgment, free riding, and production blocking, which can limit the effectiveness of the ideation process shaer2024aiaugmented. Peer judgment refers to the influence of group members’ opinions on individual contributions, which can discourage the sharing of unconventional ideas. Free riding occurs when some group members rely on others to contribute, leading to reduced individual participation. Production blocking happens when group members wait for their turn to share ideas, resulting in lost opportunities for spontaneous ideation. Online visual workspaces and the integration of LLMs into these platforms offer new avenues for enhancing group creativity by providing diverse perspectives and reducing evaluation apprehension Desmond2024. While holistic measures can be efficient, they often conflate multiple constructs, leading to inconsistencies in ratings. Therefore, a more granular approach is necessary to capture the specific aspects of idea quality systematically. For example, using a well-defined set of criteria can provide a structured and consistent evaluation framework.
Evaluating the quality of ideas is a critical aspect of the creative process. Traditional methods for idea evaluation can be broadly categorized into subjective and objective evaluations.
Subjective evaluations have traditionally been the cornerstone of idea assessment. These evaluations often involve human judges who rate the quality of ideas based on personal judgment, which can be influenced by individual biases and inconsistencies (Ben_2010). While subjective evaluations can capture nuanced insights, they are inherently limited by their lack of scalability and potential for variability Dean2006. For example, one rater may intuitively include novelty or workability in their evaluation, while another may not, leading to different ratings. Moreover, a single rater may be inconsistent across ideas because different constructs may seem more important to some ideas than others (Boudier_2023). Thus, despite their efficiency, holistic measures do not address specific evaluation components in a predictable way.
Several researchers (Fiorineschi_Rotini_2023; Kim_Maher_2023; Christensen_Ball_2016; Nelson_Wilson_Rosen_Yen_2009; Linsey_Clauss_Kurtoglu_Murphy_Wood_Markman_2011; Bryant_2005; PUCCIO2012189; Kurtoglu_Campbell_Linsey_2009; Karimi_Maher_Davis_Grace_2019; Han_Shi_Chen_Childs_2018) have reported multiple dimensions for the subjective evaluation of ideas, each capturing a specific aspect of idea quality. The following are some of the key dimensional metrics identified from various literature.
  • Novelty: The degree to which the idea is new and not derivative.
  • Rarity: The uniqueness of the idea compared to existing ideas.
  • Originality: The novelty and innovativeness of the idea.
  • Usefulness: The practical applicability of the idea.
  • Relevance: The extent to which the idea is connected to the problem statement.
  • Workability: The feasibility and implementability of the idea.
  • Diversity: The range and variety of different ideas.
  • Quantity: The total number of ideas generated.
  • Quality: The overall excellence of the idea.
  • Creativity: The degree of imagination and inventiveness in the idea.
  • Thoroughness (Specificity): The level of detail and specificity in the idea.
  • Feasibility: The practicality and achievability of the idea.
  • Effectiveness: The potential impact and success of the idea.
  • Innovativeness: The degree of innovation and breakthrough thinking in the idea.
  • Impact: The potential effect and influence of the idea.
  • Utility: The usefulness and benefits of the idea.
  • Practicality: The ease of implementation and practicality of the idea.
  • Insightfulness: The depth of understanding and insight reflected in the idea.
Objective evaluations aim to mitigate the limitations of subjective assessments by employing systematic and quantifiable metrics. Traditional automatic metrics like BLEU, ROUGE, and METEOR have been used to evaluate natural language generation systems (blagec2022globalanalysismetricsused), but these metrics often show a low correlation with human judgments (liu2023geval). To the best of the author’s knowledge, there are no objective evaluation metrics similar to subjective evaluation metrics that align with human judgment. Hence, one of the novel contributions of this research work involves developing certain dimensional metrics that correlate with human judgement.
The use of AI for evaluating ideas promises increased speed and objectivity. Various types of AI evaluation systems have been explored by several researchers, each offering unique advantages and challenges:
1. N-gram-Based Metrics: These metrics evaluate text based on term frequency and inverse document frequency. While they are simple and computationally efficient, they often fail to capture semantic meaning and contextual relevance (kim-n-gram).
2. Embedding-Based Metrics: These metrics measure the semantic similarity between generated text and reference text based on word or sentence embeddings. Examples include WMD, BERTScore, and MoverScore. Embedding-based metrics offer a more robust measure of semantic similarity by capturing contextual nuances (liu2023geval).
3. Task-Specific Evaluators: These evaluators are fine-tuned for specific tasks, such as dialogue generation or report generation. They provide tailored evaluations based on task-specific criteria (dahlander-task-specific-evaluators).
4. Unified Evaluators: These evaluators offer a generalized framework for evaluating multiple types of NLG tasks. They provide a consistent evaluation approach across different tasks and domains (zhong2022unifiedmultidimensionalevaluatortext).
5. LLM-Based Evaluators: These evaluators, such as Prometheus, use large language models to assess generative outputs based on custom score rubrics. They have shown a high correlation with human judgments and offer a scalable alternative to human evaluation. However, such a system has not been explored for idea evaluation in the product design domain. (Desmond2024).
Despite the promise of AI-based evaluation systems, challenges like prompt sensitivity, bias, and hallucinations remain. Ensuring robust and transparent evaluation processes, integrating human oversight, and validating evaluation criteria are essential for enhancing the reliability and validity of AI-based evaluations.
The rapid advancements in artificial intelligence (AI) and natural language processing (NLP) have ushered in a new era of innovation, particularly in the realm of idea generation. Conversational AI (CAI) systems, powered by large language models (LLMs) such as GPT, have been shown to demonstrate unprecedented capabilities in generating numerous ideas within a short period. This transformative potential of CAI systems necessitates the development of automated, objective methods for evaluating the vast quantity of ideas they produce.
In traditional creative processes, human designers generate ideas through brainstorming sessions, individual reflection, and collaborative discussions. The volume of ideas produced in these settings is typically manageable, allowing human experts to evaluate and rate each idea subjectively. This manual evaluation process, although time-consuming, benefits from the nuanced judgment and contextual understanding of human evaluators. However, the advent of CAI systems has fundamentally altered the landscape of idea generation. These systems can produce an overwhelming number of ideas in a fraction of the time it would take human designers. While this capability significantly enhances the creative process by providing a rich pool of potential solutions, it also introduces the challenge of idea abundance. The sheer volume of ideas generated by CAI systems renders manual evaluation impractical and inefficient.
Given the impracticality of manual evaluation, there is a pressing need for automated methods to assess and evaluate the ideas generated by CAI systems. Such methods must be objective to ensure consistency, scalability, and reliability in the evaluation process. Objective evaluation is particularly crucial in scenarios where the ideas generated are used to inform critical decision-making processes, such as product design, strategic planning, and innovation management.
This paper proposes a systematic mathematical framework for the objective evaluation of the ideation process. We can apply quantitative metrics to assess their quality, relevance, and diversity by representing ideas as mathematical entities, such as vectors in high-dimensional space. This approach enables the systematic evaluation of ideas based on clearly defined measures, eliminating the subjectivity and variability inherent in manual evaluations.
Embedding-based methods are inherently statistical, necessitating a large set of sample ideas to ensure robust and meaningful analysis. However, obtaining such extensive datasets from human participants often results in a loss of uniformity and a limited number of ideas, complicating systematic analysis. To address this challenge, we leveraged a CAI-based ideation tool, as previously published in our work, which facilitates the generation of a large set of ideas within a specific structured framework. This tool enabled us to generate the necessary sample ideas for the development and validation of our evaluation method. For the sake of completeness, the following section briefly describes the CAI-based generation process, providing context for its role in our study.
2. Idea Generation using CAI
Idea generation is a cognitively expensive process that requires experience to master (Sankar2023). Novice designers often face the struggle to generate a large variety of novel ideas. Conversational AI (CAI) systems powered by advanced natural language processing (NLP) models have been reported in our earlier work (Sankar2024) for generating a required number of ideas quickly. The conversational AI system leverages large language models to generate text based on given prompts. These models are trained on extensive datasets, enabling them to understand and generate human-like text across various domains. In the context of idea generation, (Sankar2024) established that CAI-based ideation facilitates the creative process by overcoming human cognitive limitations such as fatigue and design fixation.
0.1 2.1. Idea Representation using AOC
A Custom-built interface was developed for generating ideas, as shown in Figure 1 with labels for different fields. The response of the CAI system is managed with a field marked as ’creativity slider’, which adjusts the parameter commonly called temperature; a higher temperature setting produces more diverse outputs, while a lower temperature setting produces focused and predictable ideas.
In order for CAI to present ideas that are similarly detailed, coherent, and easily interpretable, an Action-Object-Context (AOC) model was developed. Here, action describes what the idea aims to achieve or perform; object refers to the entity involved in the action, and context provides the situation where the action happens. The ideas generated are stored as a JSON file for further investigation.
Figure 1: Interface of the Design Chatbot of the Custom-Built Conversational AI (CAI) Tool
0.2 2.2. Generating Ideas Using AOC
The problem statements serve as prompts for the CAI system, guiding the generation of relevant ideas. For this study, we used the following six distinct problem statements.
  • PS1: Product for segregation as a means for effective waste management
  • PS2: Product for footwear disinfection and cleaning for improved hygiene and safety
  • PS3: Product for enhancing household dish cleaning efficiency and sustainability
  • PS4: Product for enhancing comfort and efficiency for prolonged standing in queues
  • PS5: Product for bird-feeding for fostering mental well-being of elderly individuals at Home
  • PS6: Product for convenient umbrella drying and storage on travel
The CAI tool uses these prompts to generate 100 ideas for each problem statement, each structured according to the AOC model. A representative sample of CAI-generated ideas is shown in Table 1.  [^1] Any human-generated ideas can also be processed in subsequent sections if they are presented as per the AOC structure.
S. No. Title of the Idea Action Object Context
Problem Statement 1: Product for segregation as a means for effective waste management
1 Smart Segregation Bins Automatically sort waste Bins Use sensors and AI to identify and segregate waste into recyclables, organics, and general waste
2 Colour-Coded Waste Bags Visually indicate waste type Waste Bags Provide different coloured bags for different types of waste to encourage proper segregation at the source
3 Segregation Education App Educate and guide users Mobile App An app that teaches waste segregation and helps track the environmental impact of proper waste management
4 - 100
Problem Statement 2: Product for convenient umbrella drying and storage on travel
1 Magnetic Quick-Dry Umbrella Case Magnetically attaches and speeds up drying Umbrella Case Useful for attaching to metal doors or lockers, particularly in public areas or small apartments
2 Solar-Powered Outdoor Umbrella Stand Uses solar energy to dry umbrellas Umbrella Stand Ideal for eco-friendly outdoor storage solutions like patios or gardens
3 Automated Umbrella Spin Dryer Spins the umbrella to shed water Spin Dryer Perfect for entrance areas, offering a quick method to reduce dripping water from freshly used umbrellas
4 - 100
Problem Statement 3: Product for footwear disinfection and cleaning for improved hygiene and safety
1 UV Sanitization Mat Disinfects using UV light Mat Placed at the entrance of homes or offices to sanitize shoes upon entry
2 Self-Cleaning Smart Shoe Cabinet Automatically cleans and disinfects Shoe Cabinet For storage of shoes with built-in disinfection technology
3 Antibacterial Shoe Spray Dispenser Releases antibacterial spray Spray Dispenser A hands-free dispenser that applies a disinfectant to shoes
4 - 100
Problem Statement 4: Product for enhancing household dish cleaning efficiency and sustainability
1 Eco-Friendly Dish Scrubber Cleans using sustainable materials Dish Scrubber Made from biodegradable materials to scrub dishes effectively while being environmentally friendly
2 Water-Saving Sink Attachment Regulates water flow Sink Attachment It limits water usage during dish cleaning without compromising on cleanliness.
3 Solar-Powered Dishwasher Operates using solar energy Dishwasher Harnesses solar power to save electricity and reduce carbon footprint for household dish cleaning.
4 - 100
Problem Statement 5: Product for enhancing comfort and efficiency for prolonged standing in queues
1 Portable Foot Roller Massage and relax feet Foot Roller Provides relief from standing fatigue
2 Queue Companion Stool Provide temporary seating Stool Allows rest without losing place in the queue
3 Mobile Queue Barrier Guide and organize queue formation Barrier Streamlines queues to prevent overcrowding
4 - 100
Problem Statement 6: Product for bird-feeding for fostering mental well-being of elderly individuals at Home
1 Smart-Connect Birdhouse Facilitates remote interaction Birdhouse Allows elderly individuals to watch and interact with birds via a smartphone app, promoting a sense of connection with nature.
2 Chirp-O-Meter Feeder Measures bird activity Bird Feeder Logs the frequency of bird visits, providing elderly users with a fun way to track and anticipate bird interactions.
3 Garden Flight Paths Guides Birds Landscape Feature Designs in the garden that attract specific species to feeding spots, offering a visually stimulating activity for the elderly.
4 - 100
Table 1: A Sample List of 100 Ideas Generated by the CAI in the AOC Structure for Each of the Six Problem Statements
0.3 2.3. Advantages of Using CAI for Idea Generation
Using the CAI tool for idea generation offers the following advantages:
Speed and Efficiency: The CAI tool can generate many ideas in a fraction of the time it would take human designers. This rapid ideation accelerates the creative process and allows for exploring a broader idea space.
Overcoming Cognitive Limitations: The CAI tool is not subject to human cognitive limitations such as fatigue, design fixation, and mental blocks. This enables a more comprehensive and unbiased exploration of the idea space.
Standardization and Consistency: The AOC model provides a standardized format for representing ideas, ensuring consistency and coherence in the generated ideas. This standardization facilitates easier assessment and comparison of ideas.
3. Characteristics of Idea Exploration
Idea exploration involves generating a wide range of ideas to solve a given problem, with the aim of identifying unique, novel, and effective solutions.
0.4 3.1. Concept of Idea Space
The concept of idea space is rooted in several theoretical frameworks in creativity research and cognitive psychology. One such framework is the Geneplore Model (hunt-earl; Ward2004-mw), which posits that creative cognition involves two main processes: generation and exploration. In the generation phase, individuals produce a variety of mental representations known as pre-inventive structures. In the exploration phase, these structures are elaborated, refined, and evaluated to produce novel and useful ideas.
Another relevant theory is the Conceptual Blending Theory (Fauconnier2003), which suggests that creative ideas arise from the combination of different distinct mental sub-spaces. By blending elements from different sub-spaces, individuals can generate new ideas that transcend the limitations of any single domain.
Thus, the idea space, also known as the solution space, represents the entire range of potential ideas that can be generated to address a specific problem. For the context of this paper, the authors define Idea Space as a multi-dimensional conceptual space where each point corresponds to a unique idea. In an ideal scenario, this space should be uniformly explored to uncover diverse and innovative solutions.
Designers navigate this idea space using their cognitive abilities, drawing on their knowledge, experience, and creativity. However, the exploration of idea space is not always uniform. Designers may face cognitive bottlenecks that hinder their ability to generate a wide range of ideas, leading to clusters of similar ideas and gaps in other space areas.
0.4.1 3.1.1. Cognitive Processes Involved in Exploration of Idea Space
Exploring the idea space involves complex cognitive processes that include divergent thinking, convergent thinking, and analogical reasoning:
Divergent Thinking: The ability to generate multiple, diverse ideas from a single starting point. Divergent thinking is crucial for expanding the breadth of the idea space.
Convergent Thinking: The ability to evaluate and refine ideas to identify the most promising solutions. Convergent thinking is essential for exploring the depth of the idea space.
Analogical Reasoning: The ability to draw parallels between different domains or contexts to generate novel ideas. Analogical reasoning facilitates the blending of different conceptual spaces.
0.4.2 3.1.2. Challenges in the Exploration of Idea Space
Despite the potential for generating innovative solutions, exploring the idea space is fraught with challenges (Sankar2024). Designers often face cognitive bottlenecks that hinder their ability to explore the idea space comprehensively. These bottlenecks are classified by the authors (Sankar2024). For sake of completion, the bottlenecks are given below:
Design Fixation: A tendency to become fixated on a particular idea, limiting the exploration of the idea space.
Cognitive Biases: Systematic patterns of deviation from rationality in judgment and decision-making, such as confirmation bias and anchoring.
Lack of Domain Knowledge and/or Experience: Insufficient knowledge or experience in a particular domain can restrict the ability to generate diverse and relevant ideas from different corners of the idea space.
Mental Block: Limited time, budget, and technological resources can constrain the exploration of the idea space.
0.5 3.2. Measures for Quantifying Exploration of Idea Space
To objectively assess the exploration of idea space, we introduce two key measures: dispersion and distribution.
0.5.1 3.2.1. Dispersion
Dispersion refers to the spread or variability of the idea points in the 2D space. It indicates how much the points are spread out from each other and from the centre of the distribution across the idea space. Dispersion can be thought of as the extent to which the ideas vary from one another. High dispersion indicates that the ideas are diverse and cover a wide range of the solution space, while low dispersion suggests that the ideas are clustered in a specific region.
Key Aspects of Dispersion:
  1. 1.
    Range: The distance between the minimum and maximum values in each dimension.
  2. 2.
    Variance and Standard Deviation: Measures of how much the points deviate from the mean.
  3. 3.
    Spread: The extent to which points within a cluster are spread out.
  4. 4.
    Outliers: Points that are far away from the majority of other points, indicating extreme values or unique ideas.
0.5.2 3.2.2. Distribution
Distribution refers to the overall arrangement or pattern of the idea points in the idea space. It indicates how the points are organized or grouped and where the points are more concentrated or sparse. It refers to the uniformity with which ideas are spread across the idea space. A uniform distribution indicates that all regions of the idea space are explored equally, while a non-uniform distribution suggests that certain areas are more densely populated with ideas than others.
Key Aspects of Distribution:
  1. 1.
    Shape: The overall form of the point cloud (e.g., normal, skewed, bimodal).
  2. 2.
    Density: Areas where points are more concentrated versus areas where points are sparse.
  3. 3.
    Clusters: Groups of points close to each other, indicating similar ideas.
  4. 4.
    Patterns: Specific arrangements or structures in the data (e.g., linear, circular).
(a) 1.High Dispersion, Uniform Distribution
(b) 2.High Dispersion, Non-uniform Distribution
(c) 3.Low Dispersion, Uniform Distribution
(d) 4.Low Dispersion, Non-uniform Distribution
Figure 2: Classification of Idea Space
0.6 3.3. Classification of Idea Space
To illustrate the concept of idea exploration, we consider four possible scenarios based on the measures of dispersion and distribution:
  1. 1.
    High Dispersion and Uniform Distribution (Figure 2(a)): This scenario represents the ideal exploration of idea space, where ideas are diverse and evenly spread across the entire solution space. It indicates a comprehensive and balanced approach to idea generation.
  2. 2.
    High Dispersion and Non-Uniform Distribution (Figure 2(b)): In this scenario, ideas are diverse but clustered in certain regions of the idea space. While there is a wide range of ideas, some areas are over-explored while others are neglected.
  3. 3.
    Low Dispersion and Uniform Distribution (Figure 2(c)): This scenario represents a limited range of ideas that are evenly spread across the idea space. While the ideas are uniformly distributed, they lack diversity and may not provide innovative solutions.
  4. 4.
    Low Dispersion and Non-Uniform Distribution (Figure 2(d)): This scenario is the least desirable, where ideas are both limited in range and clustered in specific regions. It indicates a narrow and imbalanced approach to idea generation.
4. Embeddings for Idea Representation
To objectively assess idea exploration, it is essential to represent ideas as points in a mathematical sense. This allows for the application of quantitative measures to evaluate the dispersion and distribution of ideas. By converting ideas into high-dimensional vectors, we can use techniques such as embeddings to capture the semantic and contextual similarities between ideas. Embedding techniques transform natural language descriptions of ideas into dense vector representations in a high-dimensional space. These vectors can then be analyzed to objectively assess the spread and uniformity of ideas within the idea space. This provides a foundation for developing automated methods to evaluate the ideas generated by Conversational AI systems, ensuring that the exploration of idea space is both comprehensive and balanced.
Embeddings is a collective name for a set of language modelling and feature learning techniques used in Natural Language Processing (NLP) and are pivotal for the functioning of Conversational AI (CAI) systems that use Large Language Models (LLMs) like GPT (Generative Pre-trained Transformers). Embeddings are dense vector representations of words, sentences, or even entire documents. These vectors capture semantic meaning and contextual relationships between different pieces of text, enabling machines to understand and generate human language more effectively. Embeddings transform textual data into numerical form, making it possible for deep learning algorithms to process and analyze text. They are particularly useful for LLMs because they provide a way to represent complex linguistic information in a compact, numerical and computationally efficient manner.
0.7 4.1. Nature of Embeddings
Embeddings work by mapping tokens, which are subwords, words or phrases, to vectors in a high-dimensional space (typically in the order of 1000s). This mapping is typically learned during the training phase of the embedding model. The goal is to position semantically and contextually similar words close to each other in this space while dissimilar words are positioned farther apart.
0.7.1 4.1.1. Semantic Encoding
The key to the effectiveness of embeddings lies in their ability to encode semantic meaning. This is achieved through training on large corpora of text, where the model learns to predict words based on their meaning. For example, in the sentences ”The cat sat on the mat” and ”The cat sat on the rug”, embeddings capture the relationship between ”mat” and ”rug” and position them nearby in the vector space due to the frequent analogous usage.
0.7.2 4.1.2. Positional Encoding
In addition to semantic meaning, modern language models like Transformers also incorporate positional encoding to understand the context in which the words are being used. This is crucial because the meaning of a word can depend on its position in a sentence. Positional encoding adds information about the position of each word in the sequence, enabling the model to understand the structure, order and context of the text better. For example, in the sentences ”The design in the dress was detailed.” and ”She has a natural talent for fashion design.”, the usage of the word design refers to two different contexts, and this relationship is captured by the embeddings to understand the context.
0.8 4.2. Mechanism of Embeddings
Tokenization is the process of splitting text into smaller units called tokens, which can be words, subwords, or characters. For example, the sentence ”Natural Language Processing” could be tokenized into [”Natural”, ”Language”, ”Processing”] or even into subword units like [”Nat”, ”ural”, ”Lang”, ”uage”, ”Processing”] depending on the tokenization strategy. The subword strategy is employed to help the model understand spelling mistakes in words while maintaining their meaning.
The dimension of an embedding vector is a critical parameter that influences the model’s performance. Common dimensions range from 50 to 4096 or even higher for language models with larger parameters. The choice of dimension depends on the complexity of the task and the computational resources available. Higher dimensions can capture more nuanced relationships but require more computational power. While individual values do not have explicit meanings, collectively, they capture various aspects of semantic and syntactic information. For instance, certain dimensions might capture gender, tense, plurality and other linguistic attributes.
0.9 4.3. Properties and Types of Embeddings
In mathematical terms, embeddings are vectors in a high-dimensional space. They exhibit properties such as linearity, where vector arithmetic can capture analogies (for example, the vector arithmetic between the vectors of ”king” - ”man” + ”woman” $\approx$ gives the vector for ”queen”). This linearity is a powerful feature that enables embeddings to generalize across different contexts.
Embeddings can be created at various levels of granularity as follows:
  1. 1.
    Word Embeddings: Represent individual words (e.g., Word2Vec, GloVe, Text-Embedding-3).
  2. 2.
    Sentence Embeddings: Represent entire sentences (e.g., Sentence-BERT).
  3. 3.
    Paragraph Embeddings: Represent longer text segments (e.g., Doc2Vec).
For static embedding models like Word2Vec, the embeddings for a given word are constant. However, in contextual models like BERT or Text-Embedding-3, embeddings can vary depending on the word’s context within a sentence. This allows for a more nuanced understanding of language.
0.10 4.4. Types of Embedding Models
There are two types of embedding models based on their ability to capture the similarities in the text. They are as follows:
0.10.1 4.4.1. Static Embedding Models
Static embedding models capture the semantic similarity in the text. Semantic similarity measures how much the meanings of two words, phrases, or sentences are alike, regardless of their specific contexts. This concept focuses on the intrinsic meaning of the words themselves. Traditional word embedding models like Word2Vec, GloVe (Global Vectors for Word Representation), and FastText produce a single, static vector for each word, which represents its overall semantic meaning. These embeddings do not change based on context. Semantic similarity is about capturing the essence of meaning. For example, ”device” and ”instrument” would have high semantic similarity because they represent the same concept.
0.10.2 4.4.2. Dynamic Embedding Models
Dynamic embedding models capture contextual similarity in addition to semantic similarity in the text. Contextual similarity refers to the similarity between words or phrases based on the contexts in which they appear. This concept is often associated with contextual embeddings, which take into account the surrounding words to generate a representation that varies depending on the context. Contextual embeddings, like those produced by models such as BERT (Bidirectional Encoder Representations from Transformers), ELMo (Embeddings from Language Models), and Text-Embedding-3 model (OpenAI), generate different embeddings for the same word depending on its context. For example, the word ”bank” will have different embeddings in ”bank of a river” and ”savings bank”. These embeddings are generated by considering surrounding words, capturing the nuances and specific meanings in different contexts.
0.11 4.5. Process of Conversion of Text to Embedding Vectors
The process of converting text into vectors using any of the aforementioned embedding models involves several steps, such as Tokenization (Splitting the text into tokens), Embedding Generation (Mapping each token to its corresponding vector), and Aggregation (Combining the vectors to form a representation of the entire text segment). In practice, an embedding layer in a neural network architecture performs these steps. During training, the model adjusts the embedding vectors to minimize a loss function, which measures the difference between the model’s predictions and the actual outcomes. The result is a set of embeddings that capture the underlying structure and meaning of the text.
0.12 4.6. Contribution to CAI Behavior
Embeddings are crucial for the behaviour of Conversational AI (CAI) models like GPT. They enable the model to understand and generate coherent and contextually appropriate responses. By capturing the semantic and contextual relationships between words, embeddings allow the model to generate text that is not only grammatically correct but also meaningful. The use of embeddings enhances the model’s ability to understand context, making it possible to generate more relevant and accurate responses. This is particularly important in conversational AI, where understanding the context of a conversation is key to providing useful and meaningful interactions.
0.13 4.7. Demonstrating the Validity of Embeddings
In this paper, we use the Text-Embedding-3 (TE3), a dynamic embedding model from OpenAI for implementation. The following examples are shown to demonstrate the power of embeddings to translate text to vectors meaningfully.
0.13.1 4.7.1. Word Embedding
To illustrate the concept of semantic similarity between words, we compared the word chair against three sets of words using their corresponding embeddings. These sets were selected based on human judgment to represent varying degrees of similarity:
  1. 1.
    Set 1 (High Similarity): seat, sofa, bench
  2. 2.
    Set 2 (Moderate Similarity): desk, table, cushion
  3. 3.
    Set 3 (Low Similarity): light, monitor, fan
Figure 3: Heat Map of Similarity Matrix for Word Embedding
Figure 4: UMAP of Word Embedding
Embeddings for all these words were generated as $13072x$ 1 dimensional vectors using the TE3 embedding model, as shown in Table 2.
Words Generated Vector Embeddings of 3072 x 1 dimension
chair [-0.014324468, -0.019203374, -0.013582874, -0.027462386, …………0.005507309, -0.011475187, 0.004418337, 0.017657736]
seat [-0.018222768, -0.011229559, -0.002593147, -0.017478982, …………0.017478982, -0.003989768, -0.023138227, 0.007466161]
sofa [-0.009118021, 0.006891137, -0.003268764, 0.028170729, …………-0.001830129, -0.017545667, -0.022681395, -0.005337789]
bench [-0.026748225, -0.000545074, -0.02579003, -0.011465845, …………0.002626913, -0.017003881, -0.017621022, -0.001066803]
desk [-0.001566614, -0.00333886, -0.017942928, -0.021080397, …………0.022861123, -0.003298582, 0.00262445, 0.00787335]
table [0.000270125, -0.030253995, -0.016318725, 0.045984801, …………0.037722155, 0.014578803, 0.004198854, 0.016700078]
cushion [0.010662668, -0.040683772, -0.003556379, 0.016528862, …………0.016934318, 0.002203057, -0.021308083, 0.006099115]
light [-0.005160935, -0.041870821, 0.01972012, 0.010678356, …………0.023139138, -0.000709932, -0.003030125, -0.013951539]
monitor [-0.007739906, -0.00146106, -0.015165323, 0.0000751, …………0.004533872, 0.009784079, -0.011574916, -0.016606728]
fan [-0.028556051, 0.00263738, -0.019429563, 0.003650993, …………0.001529363, 0.016662996, -0.016424499, -0.006236698]
Table 2: Vectors Generated for Words using the TE3 Embedding Model
The cosine similarity between these vectors was calculated using the formula given below.
cosine $\text{cosine similarity}=\cos(\theta)=\frac{\mathbf{A}\cdot\mathbf{B}}{|\mathbf{A}||\mathbf{B}|}=\frac{\sum_{i=1}^{n}A_{i}B_{i}}{\sqrt{\sum_{i=1}^{n}A_{i}^{2}}\sqrt{\sum_{i=1}^{n}B_{i}^{2}}}$
and subsequently normalized to form a similarity matrix. The resulting heat map of this matrix is depicted in Figure 3. The heat map clearly indicates that the word ”chair” exhibits similarity scores in the range of 0.3 to 0.45 with the words in Set 1 compared to 0. to 0.25 for Set 2 and 0 to 0.1 for Set 3. As seen, the similarity scores are highest for Set 1, moderate for Set 2, and lowest for Set 3. This aligns with human judgment and demonstrates the effectiveness of embeddings in capturing semantic similarity.
To further analyze the spatial relationships between these word embeddings, a Uniform Manifold Approximation and Projection (UMAP) was calculated for all the embeddings to reduce the dimensionality from 3072 to 2. The resulting scatter plot is shown in Figure 4. The scatter plot reveals that the embeddings of the words in Set 1 are positioned closer to the word ”chair” than those in Set 2 and Set 3. This proximity indicates that the vectors for Set 1 words lie closer to the ”chair” vector, followed by Set 2 and then Set 3.
These results conclusively demonstrate that embeddings effectively capture semantic similarity between linguistic terms. The analysis shows that words with higher semantic similarity to ”chair” are represented by vectors that are closer in the high-dimensional space, validating the robustness of embeddings in representing semantic relationships in the natural language (English language in the current context).
Figure 5: Heat Map of Similarity Matrix for Sentence Embedding
Figure 6: UMAP of Sentence Embedding
0.13.2 4.7.2. Sentence Embedding
To illustrate the concept of contextual similarity, we compared the sentence Sentence 0: ”The chair was designed with ergonomic features to ensure user comfort.” against three other sentences using their corresponding embeddings. These sentences were selected to represent varying degrees of contextual similarity:
  1. 1.
    Sentence 1 (High Similarity): ”The seat incorporates ergonomic principles to maximize comfort.”
  2. 2.
    Sentence 2 (Moderate Similarity): ”The desk was built to provide ample workspace and adjustable height.”
  3. 3.
    Sentence 3 (Low Similarity): ”The light uses LED technology to provide energy-efficient lighting.”
Embeddings for these sentences were generated as $13072x$ 1 dimensional vectors using the TE3 embedding model, as shown in Table 3.
Sentences Generated Vector Embeddings of 3072 x 1 dimension
The chair was designed with ergonomic features to ensure user comfort. [-0.019350942, -0.00016339, -0.021010781, -0.025575334, …………0.010235666, -0.00643533, -0.000910317, 0.013569174]
The seat incorporates ergonomic principles to maximize comfort. [0.005157992, -0.021131463, -0.018314859, -0.032051004, …………0.017260367, -0.02217208, 0.002526965, 0.02342082]
The desk was built to provide ample workspace and adjustable height. [-0.018921912, 0.0000983, -0.010809345, -0.006264086, …………0.006312243, -0.01546944, 0.005260202, 0.000995549]
The light uses LED technology to provide energy-efficient lighting. [-0.011437991, -0.004313813, -0.013822543, -0.017071823, …………-0.013626014, 0.014582455, -0.023845525, 0.023360753]
Table 3: Vectors Generated for Sentences using the TE3 Embedding Model
The normalized cosine similarity between these vectors was calculated as mentioned above. The resulting heat map of this matrix is depicted in Figure 5.
The heat map clearly indicates that the sentence ”The chair was designed with ergonomic features to ensure user comfort.” exhibits a similarity score of 0.6 with the sentence in the high similarity category compared to 0.3 in the moderate and 0.02 in the low similarity categories.
The scatter plot of UMAP, as shown in Figure 6, reveals that the embeddings of the sentences with high contextual similarity are positioned closer to the reference sentence. The sentences with moderate and low contextual similarity are progressively farther away.
These results conclusively demonstrate that sentence embeddings effectively capture contextual similarity between sentences. The analysis shows that sentences with higher contextual similarity to the original sentence are represented by vectors that are closer in the high-dimensional space, validating the robustness of embeddings in representing contextual relationships in the natural language (English language in the present context).
5. Requirements for Idea Embedding
It is demonstrated above that embeddings reasonably capture the semantic and contextual aspects of a statement. However, it is required to establish their utility for aiding designers and the design process. It can be appreciated that the task of sifting through a large volume of ideas generated during the ideation phase is challenging even for experts to manually identify and select a few novel and diverse ideas for further development. Also, assessing the effectiveness of the idea-generation phase helps in producing fruitful and innovative solutions. Therefore, a robust and objective framework to assess the result of ideation activity would be useful. To achieve this, we aim to address the following research questions.
Semantic Validity of Idea Embedding
The Idea Embeddings are nothing but vector embeddings generated by a CAI system for the idea statements. However, the validity of these embeddings needs to be established in the design community. The following two research questions are framed to address this.
RQ1.1: Meaningfulness: is idea embedding semantically acceptable to the designers?
It is important to determine if these embeddings align with the designers’ understanding and interpretation of the ideas. By confirming the semantic validity to be in line with the designer’s expert opinion, it can be ensured that the quantitative metrics applied to these vectors reflect meaningful dimensions of the idea space, thus facilitating a more accurate and effective evaluation process.
RQ1.2: Usefulness: do idea embeddings help designers in selecting diverse ideas from a large pool of available ideas?
Although it is desirable to have a large pool of ideas to improve the chance of a good/out-of-the-box solution, it is not possible to develop all the ideas towards a practical design. It is, therefore, important to efficiently shortlist diverse and innovative ideas from the given pool, preferably without imposing a significant cognitive burden on the designers. We believe that idea embeddings can be used to objectively assess and select ideas based on their disposition in the idea space.
Characterization of Idea Space
The goal of the ideation process is to generate innovative and diverse solutions. It is necessary to evaluate how well the idea space has been explored. A uniform and comprehensive exploration ensures that no potential solutions are overlooked. The following two research questions are framed to address the need to assess the thoroughness of the idea-generation process.
RQ2.1: Distribution: Can the effectiveness of idea exploration be assessed objectively?
This question assesses whether the ideas generated are uniformly distributed over the idea space. Uniform exploration is desirable as it would ensure that no region of the idea space is over-represented or under-represented. An over-representation can be a consequence of cognitive bottlenecks experienced by designers leading to similar ideas and would manifest as clusters of ideas in the idea space.
RQ2.2: Dispersion: Can the comprehensiveness of the idea exploration be assessed objectively?
This question investigates the extent of the idea space that was covered during ideation. A high dispersion indicates high diversity and comprehensiveness of the exploration, which is desirable as it would imply that the generated ideas spanned the various dimensions of the idea space effectively.
1 6. Empirical Exploration of Idea Embedding
To answer the above questions, a CAI-based method for the generation, representation, and analysis of ideas is adopted. The following paragraphs present the steps involved in this approach.
1.1 6.1. Methodology
[…]
We propose to use techniques such as UMAP (Uniform Manifold Approximation and Projection) and DBSCAN (Density-Based Spatial Clustering of Applications with Noise) for natural clustering of points in the idea space. UMAP is chosen for its ability to preserve both local and global proximity of the data, ensuring that the intrinsic structure of the idea space is maintained even after dimensionality reduction to 2D for ease of visualization. This ensures an accurate representation of their semantic and contextual similarities in a lower-dimensional space as well.
DBSCAN clusters points (ideas in the idea space) based on their proximity and density. This technique was chosen as it effectively constructs clusters of varying shapes and sizes; it has the additional benefit of being robust to noise and outliers. Since the proximity of points or ideas can be interpreted as semantic similarity, clusters here are interpreted as groups of similar ideas.
The clusters output by DBSCAN are demarcated using dotted lines (Figure 7). It is interesting to note that the result significantly matches the perceptual organization of the points.
(a) Idea Set 1
(b) Idea Set 2
(c) Idea Set 3
(d) Idea Set 4
(e) Idea Set 5
(f) Idea Set 6
Figure 7: Cluster and Triad Plot of Vector Embeddings for Ideas Generated for the Six Problem Statements using DBSCAN and UMAP Algorithm
 [^2] R1, R2 - Reference Idea Statement 1 & 2; A - Idea from the same cluster; B - Idea from the neighbouring cluster; C - Idea from far away cluster [^3] VS - Very Similar; SS - Somewhat Similar; VD - Very Different
444# 1 to 9 - Cluster Index
1.2 6.2. Meaningfulness of Idea Embeddings (RQ1.1)
To evaluate the meaningfulness of idea embeddings, a structured experiment was conducted. From each of the six idea sets, as shown in Figure 7, two ideas were randomly selected to serve as reference ideas (R1, R2); three additional ideas were selected for each of R1 and R2 as follows. Idea-A from the cluster of R1/R2, idea-B from a neighbouring cluster and idea-C from a distant cluster. This resulted in 12 groups, each comprising one reference idea and three additional ideas, as shown in Table 1.2.1. These groups were presented in a Google Form questionnaire wherein their cluster-based similarity levels were not disclosed to the respondents.
The questionnaire was then distributed to 30 expert designers, each with more than five years of experience in the design field. The experts were asked to assess idea-A, -B and -C in relation to the reference idea and indicate them as “Very Similar (VS)”, “Somewhat Similar (SS)”, or “Very Different (VD)”. The purpose of this evaluation was to determine whether the semantic similarities and differences captured by embeddings (clusters) aligned with the judgment of human experts.
1.2.1 6.2.1. Results and Inference
The responses of the 30 experts are shown using bar plots in Figure 8. For each of the A, B and C, the dominant response (one with the highest vote) is highlighted with shading. The order of bars for each comparison idea is maintained as VS, SS, and VD from left to right.
To further elucidate the findings, the dominant responses identified by expert designers were marked on Figure 7 using arrows. These arrows originate from the reference ideas (R1 and R2), point towards each idea (A, B, and C) and are labelled as VS, SS and VD as per the dominant response. It can be observed that the dominant categorization by experts aligned well with the cluster proximity except for the cases—R1 in Idea Set 2 and R2 in Idea Set 4—where the categories VS and SS were interchanged, the VD category still aligned with the expert judgment. This consistency indicates that the idea embeddings not only capture the semantic meaning of the ideas but also preserve the similarity relationships among them, falling within the human judgment. Thus, the answer for RQ1.1 is Yes, the idea embedding is semantically acceptable to the designers.
(a) Idea Set 1
(b) Idea Set 2
(c) Idea Set 3
(d) Idea Set 4
(e) Idea Set 5
(f) Idea Set 6
Figure 8: Bar Plot of Expert Responses from Questionnaire for Similarity of Ideas Generated for the Six Problem Statements
 [^5] IS# - Idea Set [1/2/3/4/5/6] [^6] R1, R2 - Reference Idea Statement 1 & 2; A - Idea very similar to R1/R2; B - Idea somewhat similar to R1/R2; C - Idea very different to R1/R2
 [^7] VS - Very Similar; SS - Somewhat Similar; VD - Very Different
{NiceTabular}
—p0.5cm—p4.5cm—p11cm— Idea Set No. Reference Idea (R) Comparison Idea (A/B/C)
1
R1: Waste Streamline Funnel System: A funnel system that attaches to bins, guiding waste into the correct compartment based on its size and shape. A: Smart Trash Compactors: Trash compactors that can detect the type of waste and compact it accordingly, reducing volume and facilitating recycling.
B: Multi-Layer Segregation Shelves: Stackable shelves with labels for different types of waste, encouraging organized segregation in homes and offices.
C: Zero Waste Starter Packs: Kits that include reusable items and information on how to start and maintain a zero-waste lifestyle.
R2: Gamified Recycling Bins: Bins designed with game-like features that reward users with points or digital tokens for correctly segregating their waste. A: Segregation-Encouraging Trash Receptacles: Design trash receptacles with engaging visuals and sounds that reward users for proper segregation.
B: Waste Segregation Reminder System: A system that reminds users to segregate their waste at the disposal point through visual cues or audio messages.
C: Waste Type Projection System: A system that projects images onto the ground or bins to indicate where different types of waste should be thrown.
2 R1: Multi-Tiered Carousel Umbrella Dryer: Suitable for commercial spaces or public areas with heavy foot traffic needing efficient umbrella turnover. A: Hydraulic Umbrella Compression Dryer: Designed for luxury hotels or high-end residential buildings looking for an efficient and delicate drying solution.
B: Expandable Umbrella Drying Tunnel: Designed for offices or residential buildings with many users and limited space.
C: Solar-Powered Outdoor Umbrella Stand: Ideal for eco-friendly outdoor storage solutions like patios or gardens.
R2: Detachable Umbrella Drying Liner: Designed for convenience and rapid drying of personal umbrellas after arriving indoors. A: Umbrella Quick-Dry Sleeve: Convenient for on-the-go individuals who need to store their umbrellas quickly after use.
B: Modular Umbrella Rack with Drainage: Targeted at public places or businesses with a high volume of umbrella usage.
C: Umbrella Drying Lamp Post: Ideal for outdoor venues, stylishly combining lighting and drying in one solution.
3 R1: Shoe Disinfection Powder: A powder formulated to absorb excess moisture inside the shoes, which also contains antimicrobial agents to prevent bacterial and fungal growth. A: Shoe Disinfection Quick-Dry Spray: A spray that not only disinfects the shoes but also contains agents that speed up the drying process.
B: Shoe Disinfection Bubble Bath: A playful bubble bath for shoes that cleans and disinfects with antibacterial soap bubbles.
C: Shoe Sanitization and Polishing Robot: A robotic device that cleans and disinfects shoes while polishing them.
R2: Footwear Disinfection Walkway: A walkway with built-in disinfection mechanisms for public spaces like malls or airports. A: Shoe Disinfection Conveyor Belt: A conveyor belt system that cleans shoes as people walk through a checkpoint or entrance.
B: Shoe Sole Disinfection Wrap: A wrap-around device that encases the shoe sole and applies a disinfectant solution for quick cleaning.
C: Natural Herbal Shoe Disinfectant: A disinfectant that uses herbal extracts with natural antibacterial and antifungal properties to cleanse shoes safely and organically.
4 R1: Smart Dishwashing Assistant: An AI-powered device that provides real-time guidance on water usage, detergent dosage, and load arrangement to optimize dishwashing efficiency. A: Dishwasher Efficiency Monitor: A device that monitors dishwasher usage and provides recommendations on improving efficiency, such as load size and detergent amount.
B: Automated Pre-soak Dispenser: Dispenses the right amount of pre-soak solution into the sink before dishwashing to loosen debris and reduce scrubbing effort.
C: Hydrophobic Dish Coating: A nanotechnology-based coating that can be applied to dishes, making them resistant to water and food residues, thus reducing the need for scrubbing and detergent use.
R2: Ultrasonic Dish Cleaner: A device that utilizes ultrasonic waves to agitate water and remove food particles from dishes without the need for scrubbing, saving time and reducing water consumption. A: Compact Dish Sterilizer: A small, energy-efficient device that uses UV light to sterilize dishes after washing, ensuring they are bacteria-free and ready for safe use.
B: Integrated Dish Pre-Rinser: A sink attachment that pre-soaks and sprays dishes to remove food particles, easing the cleaning process and lowering water use.
C: Dishwashing Gloves with Integrated Scrubbers: A pair of dishwashing gloves with built-in scrubbing bristles, allowing for efficient cleaning without additional sponges or brushes.
5 R1: Pressure-Relief Insoles: Provides comfort to feet during long periods of standing. A: Personalized Comfort Footpad: A footpad with memory foam or smart materials that molds to the user’s feet for maximum comfort.
B: Portable Foot Roller: Relieves standing fatigue.
C: Retractable Seating Brace: A wearable device that extends to form a temporary seat.
Continued in Next Page
Table 4: Shortlisted Ideas for Questionnaire on Evaluation of Similarity
{NiceTabular}
—p0.5cm—p4.5cm—p11cm— Continued from Previous Page
Idea Set No. Reference Idea (R) Comparison Idea (A/B/C)
5 R2: Queue Wait Optimizer: A system that uses data analytics to advise users on the best times to queue up, reducing wait times. A: Queue Space Optimizer: A system that adjusts queue layouts for optimal space usage and flow.
B: Queue Space Marker: A portable device that projects a boundary on the floor to indicate personal space in crowded lines.
C: Ergonomic Queue Flooring: Ergonomically designed flooring that reduces stress on legs and back.
6 R1: Tranquil Trek Trail: A series of bird feeders placed along a garden path to encourage gentle exercise and bird watching, promoting physical and mental health. A: Garden Flight Paths: Designs in the garden that attract specific species to feeding spots, offering a visually stimulating activity for the elderly.
B: Sculpture Garden Sphere: A bird feeder that doubles as a garden sculpture, with aesthetically pleasing designs that change with user interaction, promoting creativity and visual enjoyment.
C: Nostalgic Nest Viewer: A birdhouse that includes a digital screen to display old photos or videos, triggering happy memories as birds come and go.
R2: Calm Canopy Companion: A bird feeder that plays calming forest sounds to create a serene atmosphere, promoting relaxation and stress relief. A: Serenity Stream Sanctuary: A bird feeder that includes a small, gentle water stream or fountain, providing a calming auditory experience and attracting birds with fresh water.
B: Harmony Habitat Lantern: A bird feeder with a built-in lantern that emits a soft, warm glow as the evening sets in, providing a sense of security and tranquillity.
C: Pastime Puzzle Perch: A bird feeder that features interchangeable puzzle pieces or brainteasers that the elderly can solve, stimulating cognitive function and providing a sense of achievement.
1.3 6.3. Usefulness of Idea Embeddings (RQ1.2)
To assess the practical utility of the perceptually clustered idea presentation in selecting a small subset of ideas from a large pool, a study was conducted involving 40 graduate design students (non-experts). The aim of this study was to observe both individual and group behaviours in the selection process, viz., how the visually clustered idea representation influences novice designers to select diverse ideas.
Multiple copies of each plot shown in Figure 7 were printed on paper and distributed equitably to the participants. The clusters in each plot were marked with different colours to indicate the similarity of ideas within each cluster. Each plot had 100 marked points representing 100 ideas. The participants were required to select 10 arbitrary points/ideas from 100 in 10 minutes. Textual descriptions of the ideas or the associated problems were not disclosed. However, they were informed that the clusters represented similar ideas. The selected ideas were marked with a pen on the paper.
1.3.1 6.3.1. Results and Inference
For each plot, the votes for each cluster were recorded; the division of this vote among the participants was also noted. For the purpose of analysis, the area of the convex hull of each cluster in each plot is also calculated.
Let us consider a plot that is used by a subset X of P participants. The number of participants selecting at least one idea from a cluster is its Selection Index (SI). If the number of clusters in a plot with SI=X is M where the total number of clusters is C, then we define Sampling Score (SS) for the plot as M/C. These data are presented in Figure 9, wherein the bars are arranged in the decreasing order of the cluster areas. The height of each bar represents the combined number of ideas selected by all participants from the respective clusters. The SI for each cluster is indicated at the top of each bar. The SS is indicated at the centre of each plot. From the plots the following observations can be made.
  1. 1.
    The bars follow a monotonous trend. This implies that the number of ideas selected from a cluster is positively influenced by the size of the cluster.
  2. 2.
    In 3 of 6 plots, SS=1; in 2 of the remaining 3 plots, SS is significantly close to 1. This indicates the participants predominantly made use of the clusters for their sampling. This strategy ensured that the selections were diverse, as they originated from different clusters.
  3. 3.
    In one of the cases (idea set 4), SS is low. It can be observed in Figure 7(d) that this set has 3 clusters with a significantly smaller number of ideas compared to other clusters in the set. These are the clusters with SI¡X. This shows that a cluster with less number of points has a lesser probability of selection. This observation holds true even in idea sets 3 and 6, which have one cluster each with a sparse population.
An additional observation was that, although the participants were allotted 10 minutes for the task, the visual clusters enabled them to select 10 ideas within 2 minutes. This observation underscores the effectiveness of perceptual organization techniques, such as UMAP and DBSCAN, in facilitating quick and efficient idea selection while ensuring diversity.
Thus, embedding and dimensionality reduction with clustering minimized the cognitive effort of the designers in selecting diverse and representative ideas and also provided valuable insight into the behaviour of the designers during the selection process. Thus, the answer for RQ1.2 is Yes, the idea embedding is useful to the designers.
(a) Idea Set 1
(b) Idea Set 2
(c) Idea Set 3
(d) Idea Set 4
(e) Idea Set 5
(f) Idea Set 6
Figure 9: Stacked Bar Plot of the Top 10 Selected Ideas from Each Idea Set
1.4 6.4. Distribution of Ideas in the Idea Space (RQ2.1)
Through the process of embedding, the idea space is geometrically a hypercube. An ideal idea exploration should cover this idea space with statistically uniform sampling. However, due to the cliche of the ”Curse of Dimensionality”, the assessment of uniformity in the dimension of the embedding is unreliable. Hence, we make use of the dimensionality-reduced cluster maps as shown in Figure 7 for this assessment. If an idea space is uniformly explored, the ideas would be evenly distributed across the entire area. This would prohibit clustering. However, any finite sampling exercise generally exhibits clustering. DBSCAN inherently guarantees a certain level of uniformity of points within a cluster. However, this parameter for different clusters could be different because of their adaptive nature. With this understanding, we measure the distribution of points within a cluster and the distribution of clusters in an idea space as follows.
An effective ideation exercise would have a set of ideas that cover a large region in the idea space. This means that the mean distance between two ideas in that space is large. We refer to this as a Sparse distribution. When the ideas are all diverse, no two ideas would belong to a cluster because each cluster represents a set of similar ideas. However, in the presence of clustering, the mean distance between the ideas in different clusters is usually different. Also, the larger the number of clusters, the greater the number of distinct ideas. The distinctness of ideas within each such cluster can be categorized based on their distribution into Sparse (Small No. of ideas over Larger Area), Moderate (Large/Small No. of ideas over Larger/Smaller Area respectively) and Dense (Large No. of ideas over Smaller Area). A sparse cluster is considered locally more effective than moderate, and similarly, a moderate is more effective than dense. We define Idea Sparsity (IS) to capture these distribution as follows:
$Idea\>Sparsity=\frac{A_{c}}{N_{i}}\>.(\exp({-\frac{A_{c}}{N_{i}}}))$
where $N_{i}$ is the Number of Ideas in a cluster and $A_{c}$ is the Area of the cluster computed as the area of the convex hull of the points in that cluster. If the idea sparsity is similar across all the clusters, we say that the idea exploration is uniform. In that case, in a spider plot (Figure 10), where each spoke corresponds to a cluster and the length of the spoke corresponds to idea sparsity, a uniform exploration would appear as a regular convex polygon. The larger the polygon, the more number of ideas are distributed uniformly over a larger area within each cluster.
Each cluster in an idea space is distributed over an area. Ideally, we desire to have clusters that are sparsely distributed from each other since they represent more distinct sets of ideas compared to clusters that are close to each other, as evidenced by the expert opinion in the section –. In order to quantify the nature of the distribution of clusters in the idea space, we define Cluster Sparsity (CS) as follows:
$Cluster\>Sparsity=1-\frac{\sum_{i=1}^{N_{c}}A_{i}}{A_{t}}$
where $N_{c}$ is the number of clusters in an idea space, $A_{i}$ is the area of individual clusters, and $A_{t}$ is the total area of the idea space computed as the area of the convex hull of all the points in that idea space. An exploration where the clusters are well separated would have high cluster sparsity (desirable), and a compact set of clusters will give low cluster sparsity (undesirable).
DBSCAN clusters potentially produce a default cluster with points that could not be included in any of the well-defined clusters; this is usually referred to as the noise cluster. This is the one which is not indicated with a bounding curve. The points in this cluster are important as the associated ideas are distinct from clusters of similar ideas. We compute Idea Sparsity for all clusters, including the noise cluster. However, we observe that (a) noise clusters do not significantly extend the region of explored ideas, and (b) the convex bounding curve for the noise cluster inevitably overlaps multiple identified clusters. Hence, for the computation of cluster sparsity, we discount the noise cluster.
In terms of cluster and idea sparsities, an effective idea exploration would have a high cluster sparsity with a regular polygon of large size for idea sparsity. In Figure 10, the brightness of the bounded polygon of idea sparsity represents the cluster sparsity.
The bounded polygon for each idea set is compared against the ideal regular polygon. The degree of deviation is calculated as the ratio of the area of the spider plot to the area of the largest regular polygon. This metric is referred to here as Distribution Score (DS). The polygon provides a visual measure, and the distribution score provides a quantitative measure of how well the ideas were distributed across the clusters. A close match to the regular polygon would indicate a uniform distribution, while significant deviations would suggest regions where the idea generation process may have been uneven. The distribution score for each idea set is given in Table 5. It is seen that 4 out of 6 sets have a score close to 0.5, while idea set 1 has a score of 0.61 and idea set 5 has a score of 0.77, indicating that the ideas are potentially distributed uniformly in each cluster. In terms of cluster sparsity, it can be seen that idea sets 1, 3 and 5 are brighter, showing that the clusters are well-separated compared to idea sets 2, 4 and 6, where the clusters are close to each other. Thus, overall, Idea Set 5 represents a uniform exploration compared to other idea sets in terms of idea and cluster sparsity. Thus, exploration in idea set 5 is considered the most effective, and idea set 2 is the least effective per the above metrics. Thus, the answer for RQ2.1 is, Yes, the cluster-analysis of embedded ideas provides a reasonable framework for objective assessment of the effectiveness of idea exploration.
(a) Idea Set 1
(b) Idea Set 2
(c) Idea Set 3
(d) Idea Set 4
(e) Idea Set 5
(f) Idea Set 6
Figure 10: Spider Plot of Idea Density and Cluster Density for Each Idea Set
Idea Set No. Area of the Regular Convex Polygon (ARP) Area of the Irregular Polygon (AIP) Distribution Score (DS) = AIP/ARP
1 1.45 0.89 0.61
2 2.35 1.07 0.46
3 2.28 1.19 0.52
4 1.89 1.03 0.55
5 1.16 0.90 0.77
6 1.82 0.99 0.55
Table 5: Distribution Score
1.5 6.5. Dispersion of Ideas in the Idea Space (RQ2.2)
Here, we assess whether ideas spanned all dimensions of the idea space effectively. Using the measures described above, comprehensive coverage of all the dimensions of the embedding cannot be assessed as it is based on a 2D map of the original embeddings. To assess the dispersion of ideas, which refers to the extent to which the ideas are spanned in the higher-dimensional space, Principal Component Analysis (PCA) was adopted. PCA provides eigenvectors that describe the direction and relative magnitude of dominant explorations. An effective exploration of all dimensions of idea space would give comparable eigenvalues of reasonably large magnitude. For the assessment of the comprehensiveness of different ideation exercises, the eigenvalues were computed using PCA for each idea set and sorted in descending order. The results are presented in Figure 11(a) and Figure 11(b). It was observed that for all the dimensions, the idea space is explored more or less uniformly except for the first few dominant eigen directions. The maximum eigenvalue observed was 300, although all eigenvalues eventually saturated in the range of 30 to 50. Hence, the comparison of exploration below uses the first 10 eigenvalues only.
Figure 11(a) presents the top 10 eigenvalues in descending order, which indicates the spread of ideas in different directions in the high dimensional space of idea embedding. Figure 11(b) shows the differences between subsequent eigenvalues. By examining this difference, the evenness of the spread is assessed. Ideally, for a highly dispersed idea space, the differences between subsequent eigenvalues should be smaller showing a flat line at an elevated position, indicating that the ideas are spread out across multiple dimensions. It is evident from Figure 11(b) that idea sets 6, 2 and 1 approximates a flat line, whereas idea sets 4 and 3 show a short jump between the first three eigenvalues indicating a longer spread in three dimensions and finally the idea set 5 show a significant jump between the first and second eigenvalue indicating a more directed exploration. Thus, the answer for RQ2.2 is, Yes, the analysis of the dispersion of embedded ideas in high dimension provides a reasonable framework for the objective assessment of comprehensiveness of idea exploration.
(a) Line Plot of Top 10 Highest Eigen Values of PCA Embedding for Ideas Generated for Six Problem Statements
(b) Line Plot of Absolute Different Between the Top 10 Highest Eigen Values
Figure 11: Line Plot of Eigen Values and their Difference of PCA of Embedding for Ideas Generated for Six Problem Statements
7. Discussion
An intriguing observation emerged when all idea sets were processed through UMAP and DBSCAN simultaneously. The resulting plot, as shown in Figure 12, revealed that each idea set formed a dense cluster island, representing a closely-knit group of 100 ideas. Even though in the context of the present study, the problem statements were not included as a part of the idea embeddings, certain idea sets were positioned closer to each other, suggesting potential similarities in addressing the challenges posed by their respective problems. Therefore, although individual idea sets exhibited diversity and clustering, each set, in turn, appeared as clusters in a larger context; the ideas in the respective sets have something (either object or action) in common, leading to closer clustering.
Similar plots can be used to elucidate how expert and novice designers explore ideas for various problems or how the ideation pattern changes over the career of a designer. Merging of problem-specific idea clusters could imply the presence of some versatile ideas that address multiple problems. We believe that such versatile ideas are more likely in the ideation by experts.
The analysis presented above used CAI-generated ideas, as given in  (Sankar2024), using a particular prompt structure and temperature settings. The methodology can be used to explore other prompt structures and temperature settings to derive optimally engineered prompts for design problems. This, in turn, will provide insight into the effect of a problem statement on the efficacy of ideation. Without an objective framework for generation and evaluation, as presented in this paper, the reliability of such comparative studies would always remain in question. The authors’ future research will address some of the above issues.
Figure 12: DBSCAN Clustering of UMAP Embeddings of Ideas Generated for All Six Problem Statements
 [^8] IS - Idea Set
8. Conclusion
This paper presented a novel technique for the objective evaluation and selection of ideas by leveraging advanced techniques such as Vector embedding, UMAP, DBSCAN, and PCA. The findings indicated that the use of embeddings and clustering techniques ensured that the semantic and contextual relationships between ideas were preserved, facilitating a more meaningful and efficient evaluation process. The clustered and visually organized representations of ideas enabled quick and effective decision-making, further streamlining the ideation phase. This objective method allows designers to focus on creativity and innovation rather than the tedious task of manually sifting through large volumes of ideas. This framework not only democratizes the evaluation process but also paves the way for more consistent and reproducible outcomes which otherwise would demand the expertise and experience of the designer. Finally, our research work highlights the use of CAI towards a unified process for the generation, evaluation, and selection of ideas while maintaining the necessary rigour and reduced intervention from designers.
[…]